# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition boxthm]Assumption

### v1

@@ B.2.1 Technical Lemmas
where $\xi_{sw}$ is defined by Equation (10). We observe that $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$. Thus, by using Lemma˜B.4 we have with probability at least $1-18e^{-t^{2}}$
@@
$\sup_{\theta\in\Theta}\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}$
Appendix C Computational Details
C.1 Implementation Details
C.1.1 Fair Sinkhorn
We implement the FairSinkhorn algorithm from scratch using Sinkhorn-Knopp matrix scaling as proposed in cuturi2013sinkhorn.
C.1.2 Penalized Optimal Transport
We leverage the Python OT (POT) library [flamary2021pot] throughout our experiments. To solve the penalized problem, we leverage the generic conditional gradient algorithm implemented in this package. We use the following parameters and refer to the package documentations for greater details.
@@ C.1.3 Cost Learning
We implement bilevel cost learning in Pytorch. The inner optimal transport problem is solved using the ot.sinkhorn method from POT, with a warm-started initialization of the dual variables, a stoping threshold of $10^{-6}$ and a maximum of $1000$ iterations. The used solver is sinkhorn if the entropic penalty is greater than $1$, otherwise we use sinkhorn_log. The Jacobian of the optimal transport plan is estimated by iterative differentiation [Pauwels2023]. We take a single optimizer step after each solve of the inner problem. We found Adam [kingma2014adam] to provide the best and most stable results over runs. We use its default parameters, except for the learning rate. In all experiments, we set the penalty $\mathscr{D}(\cdot,\cdot)$ to be the squared Frobenius norm between the original and parametrized cost matrix, divided by the maximum entry of the euclidean cost matrix.
C.2 Experimental Details
To generate results in Figure 3, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $2$ subgroups in each distribution. For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$. The learning rate is set to $0.1$ for the Mahalanobis and ResNet cost models, and to $0.05$ for the MLP cost model. The MLP model has $3$ hidden layers and a output dimension of $8$. The dimension of the intermediate states is set to $4$. The same architectural parameters are used for the ResNet model. We use a SiLU activation for both models. All cost learning models are trained for maximum $300$ iterations of the outer problem. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,20)$ and set $\varepsilon=2$.
To generate results in Figure 5, we generate $n_{x}=n_{y}=20$ samples for each subgroup of the $3$ subgroups in each distribution. For cost learning, we use a penalty grid of $\texttt{logspace}(-3,2,30)$ and set $\varepsilon=2$. The learning rate is set to $0.1$ for all cost learning models. The MLP model has $3$ hidden layers and a output dimension of $8$. The dimension of the intermediate states is set to $16$. The same architectural parameters are used for the ResNet model. We use a SiLU activation for both models. All cost learning models are trained for maximum $300$ iterations of the outer problem. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(-3,3,20)$ and set $\varepsilon=2$.
@@ C.2 Experimental Details
To generate results in Figure 7, we set a fixed penalty of $0.1$, a fixed entropic penalty of $\varepsilon=2$ and a fixed learning rate of $0.01$. The Mahalanobis and ResNet cost model are used with identical parameters to Figure 3. We learn a cost model on the same dataset than for Figure 3 but with $1000$ samples per subgroup. We then evaluate the fairness of each cost model on small subsamples of $50$ samples per subgroup. The results are computed for $10$ runs. For penalized OT, we tune the penalty value to $15$ to obtain similar levels of fairness, and use the same entropic penalty $\varepsilon=2$.

### v2

@@ C.2.1 Technical Lemmas
where $\xi_{sw}$ is defined by Equation (10). We observe that $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$. Thus, by using Lemma˜C.4 we have with probability at least $1-18e^{-t^{2}}$
@@
$\sup_{\theta\in\Theta}\mathbb{E}[\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\mathcal{L}_{\mathbf{F}}(\pi^{*}_{\epsilon}(c_{\theta}))\rvert]\lesssim\frac{\exp(5R_{\Theta})\epsilon^{-1}}{\sqrt{n}}.$
Appendix D Computational Details
D.1 Implementation Details
D.1.1 Fair Sinkhorn
We implement the FairSinkhorn algorithm by adapting the Sinkhorn-Knopp [Sinkhorn1967] implementation of the Python OT (POT) library [flamary2021pot, flamary2024pot].
D.1.2 Penalized Optimal Transport
We leverage the POT library [flamary2021pot, flamary2024pot] throughout our experiments. To solve the penalized problem, we leverage the generic conditional gradient algorithm implemented in this package. We use the following parameters and refer to the package documentations for greater details.
@@ D.1.3 Cost Learning
We implement bilevel cost learning in Pytorch [Paszke2019]. The inner optimal transport problem is solved using the ot.sinkhorn method from POT, with a warm-started initialization of the dual variables, a stoping threshold of $10^{-6}$ and a maximum of $1000$ iterations. The used solver is sinkhorn if the entropic penalty is greater than $1$, otherwise we use sinkhorn_log. The Jacobian of the optimal transport plan is estimated by iterative differentiation [Pauwels2023]. We take a single optimizer step after each solve of the inner problem. We found Adam [kingma2014adam] to provide the best and most stable results over runs. We use its default parameters, except for the learning rate. In all experiments, we set the penalty $\mathscr{D}(\cdot,\cdot)$ to be the squared Frobenius norm between the original and parametrized cost matrix.
The Mahalanobis matrix is initialized to the identity, while the MLP is pretrained with fairness penalty $\lambda=0$ to initialize the model close to the reference cost.
D.2 Experimental Details
In all our experiments involving MLP cost, the MLP model has $2$ hidden layers, an output dimension of $2$ and uses ReLU activation. Moreover, unless otherwise specified, the entropic regularization parameter $\varepsilon$ is set to 1.
To generate results in the top part of Figure 3 b., we generate 250 samples from the group $X$ and 25 from the group $Y$. Half of the sample has sensitive attribute 0, the other half has sensitive attribute $1$. For cost learning with Mahalanobis cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.1$. For cost learning with MLP cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.05$. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,80)$. For the vanilla OT results, we use an entropic grid of $\texttt{logspace}(0,2,20)$.
To generate results in the bottom part of Figure 3 b., we generate 250 samples from the group $X$ and 25 from the group $Y$. Half of the sample has sensitive attribute 0, the other half has sensitive attribute $1$. For cost learning with Mahalanobis cost, we use a penalty grid of $\texttt{logspace}(1,3,80)$ and the learning rate to $0.05$. For cost learning with MLP cost, we use a penalty grid of $\texttt{logspace}(0,4,80)$ and the learning rate to $0.01$. For the penalized OT results, we use a penalty grid of $\texttt{logspace}(0,3,80)$. For the vanilla OT results, we use an entropic grid of $\texttt{logspace}(0,2,20)$.
@@ D.2 Experimental Details
To generate results in Figure 4, we set a fixed entropic penalty of $\varepsilon=1$. For the penalized approach, the fairness penalty is set to 90. The the cost learning approach with Mahalanobis, the fairness penalty is 1000 and the learning rate is 0.1 while for the MLP, the fairness penalty is 500 and the learning rate is 0.05. The cost are learned using the mixture of Gaussian dataset with 1000 samples from $x$ and 100 samples from $y$. We then evaluate the fairness of each cost model on subsamples of $500$ samples from $x$ and $500$ samples from $y$. The results are computed for $10$ runs.